If a bike shop determines that when x hundred bikes are built, the average cost per bike is given by c(x)=0.6^2-0.8x+5.854 where c(x) is in hundreds dollars. How many bikes should be built to mininize the average cost per bike.
c(x)=0.6x^2-0.8x+5.854
c(x)=0.6(x^2- ?)+5.854
Thank you!

Answer by stanbon(75887)

If a bike shop determines that when x hundred bikes are built, the average cost per bike is given by c(x)=0.6^2-0.8x+5.854 where c(x) is in hundreds dollars. How many bikes should be built to mininize the average cost per bike.
c(x)=0.6x^2-0.8x+5.854
---------------
Cost is a quadratic with a = 0.6, b = -0.8
-----
Minimum cost occurs when x = -b/2a = 0.8/(2*0.6) = 0.2/0.3 = 0.6667
-------
# of bikes should be 0.667*100 = 66 or 67
=============
Cheers,
Stan H.